when nitrogen gas reacts with hydrogen gas,ammonia gas is formed.How many grams of hydrogen gas are required to react completely with 3,5 L of nitrogen at STP?
1
Expert's answer
2020-11-18T13:43:29-0500

Solution:

A chemical equation must be balanced:

N2 + 3H2 = 2NH3

According to the reaction equation for 3 mol of H2 is needed 1 moles of N2.

Therefore,

mass of H2 = V of N2 / 22,4L/mol × 3mol H2 / 1mol N2 × M of H2 = 3.5L N2/22.4L/mol × 3mol H2/1mol N2 ×2 g/mol H2 = 0.9375 g H2

Answer: mass of H2 is 0.9375 g